# xApp Conflict Mitigation with Scheduler

### v2

@@ Abstract
Open RAN (O-RAN) fosters multi-vendor interoperability and data-driven control but simultaneously introduces the challenge of coordinating pre-trained xApps that may produce conflicting actions. Although O-RAN specifications mandate offline training and validation to prevent the deployment of untrained or inadequately tested models, operational conflicts can still arise under dynamic and context-dependent conditions.This work proposes a scheduler-based conflict mitigation framework to address these challenges without requiring training xApps together or further xApp re-training. By examining an indirect conflict involving power and resource block allocation xApps and employing an Advantage Actor-Critic (A2C) approach to train both xApps and the scheduler, we illustrate that a straightforward A2C-based scheduler improves performance relative to independently deployed xApps and conflicting cases. Notably, among all tested deployment scenarios (including individual xApp deployment, multiple conflicting xApps, and limited scheduler configurations), augmenting the system with baseline xApps and enabling the scheduler to select from a broader pool achieves the highest total transmission rate, thereby underscoring the importance of adaptive scheduling mechanisms. These findings highlight the context-dependent nature of conflicts in automated network management, as two xApps may conflict under certain conditions but coexist under others. Consequently, the ability to dynamically update and adapt the scheduler to accommodate diverse operational intents is vital for future network deployments. By offering dynamic scheduling without re-training xApps, this framework advances practical conflict resolution solutions while supporting real-world scalability.
{IEEEkeywords}
@@
1 INTRODUCTION
The Radio Access Network (RAN)—a critical and costly component of cellular systems—is pivotal to 6G innovation [1] and shifting from monolithic and hardware-centric architectures to more agile and customizable frameworks [2]. The Open RAN (O-RAN) paradigm, widely regarded as a cornerstone of future 6G networks [3], accelerates this transition by disaggregating traditional RAN architectures and standardizing interfaces among key network functions, thereby increasing vendor flexibility while reducing deployment and maintenance costs [4]. By interconnecting virtualized components via open interfaces, O-RAN enables multi-vendor interoperability and data-driven closed-loop control, facilitating real-time monitoring and more sophisticated RAN interactions that culminate in flexible, cost-effective, and vendor-neutral network deployments.
@@ 1 INTRODUCTION
The O-RAN Alliance adopted the $7.2$x functional split to reconcile the simplicity of the O-RU with the data rate and latency constraints at the O-RU/O-DU interface (i.e. Open Fronthaul (FH)) [9] and adopted split option $2$ between the O-DU/O-CU [10, 11]. The O-CU manages higher-layer non-real-time processes (e.g., packet management) by serving as a logical node to host higher protocol layers—namely RRC, SDAP, and PDCP [12]. It is functionally divided into the control plane (O-CU-CP), which handles RRC procedures and the PDCP control segment, and the user plane (O-CU-UP), which manages SDAP and the PDCP user segment. These components interact via the E1 interface, thereby enabling flexible scaling and reduced deployment costs for user-plane processing [4, 13]. The O-DU is also a logical node to oversee time-sensitive lower-layer operations and is responsible for hosting the lower-layer protocols of the RAN stack; specifically, the high PHY, MAC, and RLC layers—thereby executing baseband processing tasks such as scrambling, modulation, partial pre-coding, and physical resource block allocation [14, 15]. It manages multiple O-RUs over the open fronthaul and provides essential functionalities for data segmentation, scheduling, and multiplexing, with the MAC layer generating transport blocks for the physical layer based on data buffered at the RLC layer [16, 9]. This disaggregated design, controlled by the O-CU, enhances deployment flexibility, enables tighter synchronization of lower-layer operations, and allows for more efficient resource utilization. Finally, the O-RU is a physical node dedicated to low PHY layer and RF processing by performing time-domain operations—such as precoding, FFT/IFFT, and cyclic prefix addition/removal—the O-RU remains cost-effective to deploy and maintain at the cell site.
Within O-RAN, as given in Fig. 1, key entities called RAN Intelligent Controllers (RICs) host intelligent applications that dynamically and efficiently reconfigure the RAN. Specifically, third-party applications—rApps in the Non-Real-Time RIC (Non-RT RIC) and xApps in the Near-Real-Time RIC (Near-RT RIC)—operate independently of the underlying vendor to manage and optimize RAN functionalities under different latency requirements [9]. The Non-RT RIC, operating at time scales longer than $1s$, complements the near-RT RIC’s intelligent RAN operation and optimization by offering guidance, enrichment information (EI), and AI/ML model management. Meanwhile, the Near-RT RIC, deployed at the network edge and operating control loops with periods ranging from $10ms$ to $1s$, hosts xApps—each a microservice specifically designed for radio resource management through specialized interfaces and service models. The Near-RT RIC can be deployed either at the edge or within the regional cloud [9, 17], whereas the Non-RT RIC is commonly hosted by the SMO, which manages the configuration and orchestration of RAN components [14, 6].
@@
2 BACKGROUND
@@
The potential conflict types among xApps were introduced under the Near-RT RIC architecture [32]. Recently, the O-RAN Alliance’s Working Group 3 (WG3) published an independent technical specification on conflict mitigation [33], analyzing these conflict types and detailing the methods for conflict detection, resolution, and avoidance. This functionality is critical, as multiple xApps—each pursuing distinct objectives—can produce conflicting actions when deployed concurrently.According to O-RAN technical specifications in  [33], these conflicts among xApps (i.e. $\mathcal{X}_{1},\mathcal{X}_{2},\ldots$) are categorized as follows:
Figure 2: Conflict Types
@@
  • Direct conflict: Occurs when multiple xApps request incompatible adjustments to the same configuration of one or more shared network control parameters (NCPs) (i.e. ${\lambda_{1},\lambda_{2},\ldots}$) within the O-RAN. As illustrated in Fig. 2(a), if two xApps, $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$, attempt to modify the same NCP, $\lambda_{2}$, in opposite directions or with inconsistent magnitudes, a conflict ensues. For example, $\mathcal{X}_{1}$ may increase transmission power to reduce cell-edge outage, while $\mathcal{X}_{2}$ may decrease it to mitigate inter-cell interference and enhance capacity. Similarly, $\mathcal{X}_{1}$ may raise antenna tilt to expand coverage, whereas $\mathcal{X}_{2}$ may lower tilt to confine signal propagation and reduce interference.
@@
  • Indirect conflict: Arises when multiple xApps modify different NCPs to optimize the same key performance metrics (KPMs) (i.e. $\mathcal{K}_{1},\mathcal{K}_{2},\ldots$), thereby introducing unintentional interference with each other’s objectives. As illustrated in Fig. 2(b), $\mathcal{X}_{3}$ adjusts $(\lambda_{4},\lambda_{5})$ while $\mathcal{X}_{4}$ adjusts $(\lambda_{6},\lambda_{7})$ — yet both aim to optimize the same KPM, $\mathcal{K}_{4}$, in opposite directions or with inconsistent target values. For instance, $\mathcal{X}_{3}$ might adjust the cell individual offset to balance cell load, while $\mathcal{X}_{4}$ modifies the cell’s electrical tilt to enhance capacity—both altering the effective cell boundary and thus creating an indirect conflict on the same KPM[34].
@@
  • Implicit conflict: Appears when multiple xApps independently optimize distinct KPMs by tuning different NCPs, yet inadvertently impact each other’s objectives. These conflicts often remain hidden because the interdependence between the xApps is not immediately evident. As illustrated in Fig. 2(c), $\mathcal{X}_{5}$ tunes $(\lambda_{8},\;\lambda_{9})$ while $\mathcal{X}_{6}$ modifies $(\lambda_{10})$, each pursuing distinct KPMs. However, $\mathcal{K}_{8}$ influences one of $\mathcal{X}_{5}$’s NCPs, thereby implicitly affecting $\mathcal{K}_{7}$. For example, an Energy Efficiency (EE) xApp may aim to minimize overall power consumption by deactivating certain RUs, while a Slice Management (SM) xApp simultaneously strives to meet high resource demands for bandwidth-intensive applications. These competing goals—reducing energy usage vs. maximizing throughput—generate an implicit conflict [35].
@@
Unless an accurate digital twin of the network is available, it is impractical to assume that KPMs can be accurately estimated for every scenario involving diverse combinations of active xApps. Moreover, given that all xApps have been trained, tested, and verified, they cannot be modified further [9, 31]. In practical conflict management scenarios, it is essential to recognize that conflicts are often context-dependent (i.e. operational conditions), meaning that two applications may conflict in certain circumstances but coexist in others. Moreover, predicting which applications will be active concurrently is inherently difficult, making it unlikely that those trained together will ultimately operate together in real-world deployments.
This study introduces a framework that incorporates a scheduler for selecting active xApps based on context variables and a specified KPM, without requiring xApp re-training. It has been shown that independently deployed xApps can lead to conflicts, which are also shown to be highly context-dependent. To address this, an intent-driven, scheduler-based conflict mitigation framework is defined, exemplified by a simulation scenario of indirect conflict involving Power and RGB allocation xApps. The Advantage Actor-Critic (A2C) method is employed to train both the xApps and the scheduler, with two distinct scheduling approaches determining which xApps are activated according to contextual variables and the target KPM specified in the user intent.
The paper is organized as follows: In Section 3, the system model is presented, including a summary spanning from the REINFORCE algorithm to A2C methods, and the modeling of power and RGB allocation xApps using A2C. In Section 4, the proposed scheduler management framework with associated RIC components is introduced. Subsequently, Section 5 provides the training results of the xApps, the proposed methods for scheduler design, and simulation results. Finally, conclusions are drawn in Section 6.
3 SYSTEM MODEL

### v3

@@ Abstract
Open RAN (O-RAN) promotes multi-vendor interoperability and enables data-driven control, yet it also gives rise to the challenge of coordinating independently pre-trained xApps whose decisions may conflict. Although the O-RAN architecture mandates offline training and validation to ensure that only rigorously tested models are deployed, dynamic network conditions can still precipitate operational conflicts. These contextual network factors exert a direct influence on conflict occurrence and must therefore be integrated into any resolution mechanism. To this end, a context-aware conflict mitigation framework based on a scheduler approach is proposed in this work, by which competing xApp actions are reconciled through adaptation to real-time network context—without requiring joint training of xApps or further retraining. Through examination of an indirect conflict involving power and resource block allocation xApps and the employment of an Advantage Actor-Critic (A2C) approach to train both the xApps and the scheduler, it is demonstrated that performance is improved by a scheduler trained with contextual variables, relative to independently deployed xApps (i.e., the conflicting case). Moreover, the system attains its highest performance by incorporating baseline xApps and granting the scheduler access to an expanded action space, thereby emphasizing the pivotal role of adaptive scheduling mechanisms. Through these findings, the context-dependent nature of conflicts in automated network management is underscored, since two xApps may conflict under specific conditions yet coexist under alternative scenarios. Consequently, the ability to dynamically update and adapt the scheduler to accommodate diverse operational intents is vital for future network deployments. By enabling dynamic scheduling without necessitating xApp re-training, the proposed framework is positioned to advance practical conflict resolution solutions while ensuring scalability.
Index Terms:
@@
I INTRODUCTION
The radio access network (RAN)—a critical and costly component of cellular systems—plays a pivotal role in advancing 6G innovation [1]. Once confined to monolithic, hardware-centric architectures, the RAN has evolved into agile and highly customizable frameworks [2], substantially reducing human intervention and bolstering operational efficiency through AI/ML-driven control loops [3, 4]. By supporting zero-touch, multi-vendor ecosystems, these developments lay the groundwork for 6G [5, 6], where automation underpins the delivery of diverse, large-scale services. At the same time, heightened autonomy involving multiple independent control loops inevitably leads to scenarios in which individual controllers, each pursuing distinct optimization objectives, may operate in conflict in the absence of effective coordination. To illustrate how such conflicts emerged and were addressed under earlier automation paradigms, Self-Organizing Networks (SON) and Intent-Driven Networking (IDN) are revisited. Building on these historical insights, the potential of Open RAN (O-RAN), a multi-vendor and data-driven architecture, is examined, highlighting its promise for future wireless ecosystems while also magnifying the risk of uncoordinated or conflicting actions among xApps, third-party applications.
One of the earliest large-scale initiatives to integrate autonomous functionalities into cellular networks was the SON, developed under the auspices of the 3GPP [7] and the next generation mobile network (NGMN) [8] alliance to address the growing complexity of 4G systems while simultaneously reducing operational expenses and enhancing service quality [9, 10, 11]. By employing automated closed-loop control processes for self-configuration, self-optimization, and self-healing, SON functions (SONFs) significantly reduce the need for manual intervention in tasks such as neighbor relation management, antenna tuning, and outage recovery [12]. This approach not only streamlines operational workflows but also sets the stage for more advanced, intelligent automation in the RAN. However, this same autonomy inherently generates potential conflicts whenever multiple SONFs—each targeting the objective of a distinct key performance metric (KPM) (e.g., coverage, capacity, or energy efficiency)—concurrently adjust overlapping network control parameters (NCPs) [13].
The conflicts among SONFs can be broadly categorized into four types: NCP, KPM, logical dependency, and measurement conflicts, each rooted in distinct mechanisms of mutual interference [14, 15, 16, 17]. Frequently observed NCP conflicts arise when multiple SONFs attempt to configure the same NCP (e.g., transmission power) in opposing directions or magnitudes [18, 19], while KPM conflicts occur when different SONFs influence the same KPM (e.g., throughput) via separate NCPs (e.g., adjusting antenna tilt or handover offsets) [20, 21]. Additionally, logical dependency conflicts emerge when the outcome of one SONF directly shapes another’s operating conditions [22], and measurement conflicts surface when decision-making relies on outdated or asynchronous data, leading to suboptimal parameter updates [17].
Early conflict-resolution approaches in SON primarily relied on policy-based strategies that apply static rules to govern SONF actions, including assigning priorities [23, 24], implementing impact-time sequencing [25], specifying predefined parameter ranges [26], or truncating parameter updates at a defined threshold [27]. Although straightforward to implement, these methods lack the flexibility to adapt to evolving network conditions. Meanwhile, co-designing unifies multiple SONFs into a single entity [28], leveraging joint optimization [20] or decision trees [29], albeit at the cost of increased complexity. More advanced objective-driven approaches align SONF behavior with high-level operator goals by computing feasible parameter sets in real time, although they must contend with the exponential growth of state spaces and interdependencies among SONFs [30, 31]. Meanwhile, closed-loop frameworks exploit configurable factors such as mutual influences and priorities to fine-tune SONF outputs, but the intricate relationships between NCPs and KPMs remain challenging to model precisely [32, 33]. In response, recent research has moved toward proactive and cognitive coordination through data-driven algorithms [34, 35], including support vector machine (SVM)-based predictions [36] and multi-agent reinforcement learning (MARL) [37, 38, 39], which enhance the system’s ability to detect and mitigate emerging conflicts under dynamic network conditions.
Building on SON developments, IDN has emerged as a transformative paradigm for 5G RAN management [40] by transitioning from rigid, low-level configurations to high-level outcome specifications that reduce operational complexity and foster agile, scalable, and efficient orchestration [41, 42, 43]. Recent efforts highlight the potential of IDN to simplify network management in 5G and beyond, including a “one-touch” multi-domain slicing platform that minimizes manual intervention [44], an intent-based orchestrator for dynamically relocating applications in cloud-native 5G networks [45], an NLP-augmented solution for slice provisioning and service deployment tailored to private 5G environments [46], and a slicing approach that leverages mixed integer programming to optimize QoS fulfillment in 6G [47]. Although IDN is intended to streamline management and autonomously coordinate multiple RAN operations, it can inadvertently engender conflicts when multiple high-level intents—each encapsulating distinct or overlapping performance objectives and imposing divergent requirements on shared network resources—are deployed concurrently. In [48], these conflicts are classified into four principal types: (i) direct parameter conflicts, which occur when separate intents prescribe contrasting adjustments to the same NCPs; (ii) dependency conflicts, arising when interdependent control loops generate circular or inconsistent logic; (iii) contextual conflicts, which manifest only under specific operational conditions (e.g., high user demand or changing mobility patterns) and remain dormant otherwise; and (iv) latent conflicts, encompassing unintended negative outcomes that are not immediately apparent but become evident only upon observing performance degradation or user-experience issues.
To address conflicts in IDN, initial game-theoretic strategies focus on optimizing parameter values via the Nash social welfare function [49], where each KPM is managed by a cognitive function, and direct conflicts are reconciled by selecting a globally efficient equilibrium. This approach is extended with a Fisher market model [50] to accommodate varying intent priorities, achieving higher performance and reduced computation time, and further refined through diverse bargaining solutions [51] to ensure Jain-based fairness. Meanwhile, penalty-based frameworks model the cost of failing or partially fulfilling each intent’s KPM as a fitness value [52]; an evaluation agent then calculates and minimizes the aggregated penalty across all active intents, executing the derived adjustments through closed-loop control. Additionally, a gradient-based approach using the multiple gradient descent algorithm (MGDA) simultaneously minimizes multiple intent-specific loss functions by deriving a unified descent direction, thus offering a systematic way to improve performance across all tasks concurrently [53]. More advanced MARL techniques [54] allocate resources based on intent-defined priorities (e.g., maximum bit rate), thereby allowing lower-priority objectives to be proportionally degraded in favor of more critical services while maintaining an automated, data-driven resolution of competing requirements.
As it is noted, this persistent challenge of conflict for both SON and IDN frameworks is that independent autonomous optimization functionalities or control loops may inherently conflict in the absence of overarching coordination. Multiple agents, each pursuing distinct yet interdependent objectives, can undermine one another’s performance by independently tuning overlapping parameters or competing for limited resources. This concern becomes even more pronounced with the advent of O-RAN architectures, wherein disaggregated network components and third-party applications (e.g., xApps, rApps). The O-RAN architecture, widely regarded as a cornerstone of future 6G networks [55], accelerates the trend toward disaggregation and open interfaces [56], enabling multi-vendor interoperability and data-driven control loops. Such openness not only promotes a more diverse ecosystem of third-party xApps but also amplifies the risk of conflict due to heterogeneous training data, optimization targets, and vendor-specific deployments. Consequently, managing these potential conflicts in a scalable, dynamic manner becomes a pressing challenge, particularly under the O-RAN specification that calls for offline training and validation but does not fully prevent operational conflicts under dynamic conditions.
In response to these complexities, this work proposes a context-aware conflict mitigation framework via a scheduler for O-RAN that obviates the need to retrain or co-train xApps explicitly for collaborative operation. Focusing on an indirect conflict scenario involving power allocation and resource block group (RBG) allocation xApps, an Advantage Actor-Critic (A2C) approach is employed to independently train both the xApps and the scheduler. Experimental results demonstrate that even a straightforward scheduler, trained on contextual variables and responsible for selecting the appropriate xApp action, substantially reduces conflicts and improves the total transmission rate compared with independently deployed, potentially conflicting xApps. Furthermore, enabling the scheduler to choose from an expanded pool with baseline xApps achieves the highest overall performance.
The remainder of this paper is organized as follows: Section II introduces a brief O-RAN structure, the types of conflicts in O-RAN, surveys existing work on conflict management, identifies key open challenges, and summarizes the motivation for this study. Section III presents the system model, including an overview of reinforcement-learning techniques from REINFORCE to A2C, and details how the power and RBG allocation xApps are modeled using A2C. Section IV then describes the proposed scheduler-management framework along with its associated RIC components. Subsequently, Section V provides training results for the xApps, outlines the scheduler-design methods, and discusses simulation outcomes. Finally, conclusions are drawn in Section VI.
II BACKGROUND
II-A Brief O-RAN Architecture
@@
The O-RAN Alliance adopted the 7.27.2x functional split to reconcile the simplicity of the O-RU with the data rate and latency constraints at the O-RU/O-DU interface (i.e. open fronthaul (FH)) [61] and adopted split option $2$ between the O-DU/O-CU [62, 63]. The O-CU manages higher-layer non-real-time processes (e.g., packet management) by serving as a logical node to host higher protocol layers, namely RRC, SDAP, and PDCP [64]. It is functionally divided into the control plane (O-CU-CP), which handles RRC procedures and the PDCP control segment, and the user plane (O-CU-UP), which manages SDAP and the PDCP user segment. These components interact via the E1 interface, thereby enabling flexible scaling and reduced deployment costs for user-plane processing [56, 65]. The O-DU is also a logical node to oversee time-sensitive lower-layer operations and is responsible for hosting the lower-layer protocols of the RAN stack; specifically, the high PHY, MAC, and RLC layers—thereby executing baseband processing tasks such as scrambling, modulation, partial pre-coding, and physical resource block allocation [66, 67]. It manages multiple O-RUs over the open fronthaul and provides essential functionalities for data segmentation, scheduling, and multiplexing, with the MAC layer generating transport blocks for the physical layer based on data buffered at the RLC layer [68, 61]. This disaggregated design, controlled by the O-CU, enhances deployment flexibility, enables tighter synchronization of lower-layer operations, and allows for more efficient resource utilization. Finally, the O-RU is a physical node dedicated to low-PHY-layer and RF processing; by handling time-domain operations, such as precoding, FFT/IFFT, and cyclic-prefix addition/removal, it remains cost-effective to deploy and maintain at the cell site.
Within O-RAN, as given in Fig. 1, key entities called RAN intelligent controllers (RICs) host intelligent applications that dynamically and efficiently reconfigure the RAN. Specifically, third-party applications, including rApps in the non-real-time RIC (Non-RT RIC) and xApps in the near-real-time RIC (Near-RT RIC), operate independently of the underlying vendor to manage and optimize RAN functionalities under varying latency requirements [61]. The Non-RT RIC, operating at time scales longer than $1s$, complements the near-RT RIC’s intelligent RAN operation and optimization by offering guidance, enrichment information (EI), and AI/ML model management. Meanwhile, the Near-RT RIC, deployed at the network edge and operating control loops with periods ranging from $10ms$ to $1s$, hosts xApps—each a microservice specifically designed for radio resource management through specialized interfaces and service models. The Near-RT RIC can be deployed either at the edge or within the regional cloud [61, 69], whereas the Non-RT RIC is commonly hosted by the SMO, which manages the configuration and orchestration of RAN components [66, 58].
@@
II-B Conflict Types in O-RAN
Figure 2: Conflict types defined by O-RAN Alliance (WG3)
@@ II-B Conflict Types in O-RAN
The potential conflict types among xApps were introduced under the Near-RT RIC architecture [84]. Recently, the O-RAN Alliance’s working group 3 (WG3) published an independent technical specification on conflict mitigation [85], analyzing these conflict types and detailing the methods for conflict detection, resolution, and avoidance. This functionality is critical, as multiple xApps, each pursuing distinct objectives, can produce conflicting actions when deployed concurrently. According to O-RAN technical specifications in  [85], these conflicts among xApps (i.e. $\mathcal{X}_{1},\mathcal{X}_{2},\ldots$) are categorized as follows:
@@
  • Direct conflict: Occurs when multiple xApps request incompatible adjustments to the same configuration of one or more shared NCPs (i.e. $\lambda_{1},\lambda_{2},\ldots$}) within the O-RAN. As illustrated in Fig. 2(a), if two xApps, $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$, attempt to modify the same NCP, $\lambda_{2}$, in opposite directions or with inconsistent magnitudes, a conflict ensues. For example, $\mathcal{X}_{1}$ may increase transmission power to reduce cell-edge outage, while $\mathcal{X}_{2}$ may decrease it to mitigate inter-cell interference and enhance capacity. Similarly, $\mathcal{X}_{1}$ may raise antenna tilt to expand coverage, whereas $\mathcal{X}_{2}$ may lower tilt to confine signal propagation and reduce interference.
@@
  • Indirect conflict: Arises when multiple xApps modify different NCPs to optimize the same KPMs (i.e. $\mathcal{K}_{1},\mathcal{K}_{2},\ldots$), thereby introducing unintentional interference with each other’s objectives. As illustrated in Fig. 2(b), $\mathcal{X}_{3}$ adjusts $(\lambda_{4},\lambda_{5})$ while $\mathcal{X}_{4}$ adjusts $(\lambda_{6},\lambda_{7})$ — yet both aim to optimize the same KPM, $\mathcal{K}_{4}$, in opposite directions or with inconsistent target values. For instance, $\mathcal{X}_{3}$ might adjust the cell individual offset to balance cell load, while $\mathcal{X}_{4}$ modifies the cell’s electrical tilt to enhance capacity—both altering the effective cell boundary and thus creating an indirect conflict on the same KPM[86].
@@
  • Implicit conflict: Appears when multiple xApps independently optimize distinct KPMs by tuning different NCPs, yet inadvertently impact each other’s objectives. These conflicts often remain hidden because the interdependence between the xApps is not immediately evident. As illustrated in Fig. 2(c), $\mathcal{X}_{5}$ tunes $(\lambda_{8},\;\lambda_{9})$ while $\mathcal{X}_{6}$ modifies $(\lambda_{10})$, each pursuing distinct KPMs. However, $\mathcal{K}_{8}$ influences one of $\mathcal{X}_{5}$’s NCPs, thereby implicitly affecting $\mathcal{K}_{7}$. For example, an Energy Efficiency (EE) xApp may aim to minimize overall power consumption by deactivating certain RUs, while a Slice Management (SM) xApp simultaneously strives to meet high resource demands for bandwidth-intensive applications. These competing goals of reducing energy usage and maximizing throughput generate an implicit conflict [87].
II-C Related Works
@@
II-D Open Challenges and Motivation
While few solutions have been proposed for detecting and resolving conflicts among xApps (as discussed above), some critical challenges remain underexplored. These challenges must be addressed to ensure robust and adaptive conflict-management mechanisms in real-world O-RAN deployments:
  • Reliance on accurate KPM estimation: It is overly optimistic to assume that KPMs can be accurately estimated for every possible combination of xApps and their respective actions across diverse network conditions. Consequently, existing prioritization-based approaches that depend on KPM estimation [87, 88, 90] remain heavily reliant on the availability of a highly accurate digital twin of the network.
  • Uncertainty in concurrent xApp deployments: In practical environments, accurately predicting which xApps will be concurrently active is inherently challenging. It is unlikely that xApps trained together offline [92, 93, 94] will ultimately operate together in real deployments, creating scenarios that deviate from the assumptions made during pre-training.
  • Immutability of deployed xApps: As stipulated in [61, 83], once xApps have been trained, tested, and verified, they cannot be modified after deployment. This immutability poses significant challenges for conflict mitigation, as operators cannot simply retrain or adapt xApp policies (e.g., exchanging actions among xApps by incorporating these actions into the state space, as in [92, 93]) to address conflicts that emerge in dynamic network environments.
  • Context-dependent nature of conflicts: In conflict-management scenarios, it is vital to acknowledge that conflicts are context-dependent [89]. Specifically, varying operational conditions (e.g., user-equipment density, traffic load patterns) can cause two xApps to conflict under certain circumstances while allowing them to coexist under others. However, many existing solutions fail to consider these contextual nuances, resulting in over-generalized or static conflict-management strategies.
These observations reveal the shortcomings of current xApp conflict-management strategies and highlight the necessity for more adaptable and context-aware solutions in practical O-RAN deployments. To this end, this work proposes a scheduler-based framework that activates xApps according to explicit contextual variables and user-defined KPM targets, thereby eliminating the requirement to retrain existing xApps. First, we demonstrate the framework through a simulation scenario of Power and RBG Allocation xApps—each trained using an A2C algorithm—which display a pronounced, context-dependent indirect conflict. To address this issue, we then introduce an intent-driven conflict-mitigation mechanism that constrains the deployable xApps space: the scheduler dynamically selects and orchestrates xApp activations in alignment with prevailing network operational context conditions and specified performance objectives. Finally, we detail two distinct scheduling strategies that exploit contextual inputs and target KPMs derived from user intent, thus offering a flexible and efficient approach to conflict mitigation within O-RAN environments.
III SYSTEM MODEL